15m+3 the whole thing over 2 = -6


3/2x-15=-12


-7x+3=-15/2


4x+1/2(12x+2)=3


3/2(x+8)=6x

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:

Hello Student,
Please find answer to your question below

\frac{15m+3}{6} = -3
15m+3 = -18
m = \frac{-21}{15} = \frac{-7}{5}
\frac{3}{2x-15} = -12
2x-15 = \frac{-1}{4}
x = \frac{-59}{8}
-7x+3 = \frac{-15}{2}
-7x = \frac{-21}{2}
x = \frac{3}{2}
\frac{4x+1}{2(12x+2)} = 3
4x+1 = 72x+12
x = \frac{-11}{68}
\frac{3}{2(x+8)} = 6x
\frac{1}{2(x+8)} = 2x
4x^{2}+32x-1=0
x = \frac{-4\pm \sqrt{32^{2}+16}}{8}
x = \frac{-4\pm 4\sqrt{65}}{8}
x = \frac{-1\pm \sqrt{65}}{2}